### v1

# From Dataset to Real-world: General 3D Object Detection via Generalized Cross-domain Few-shot Learning

Abstract
LiDAR-based 3D object detection datasets have been pivotal for autonomous driving, yet they cover a limited range of objects, restricting the model’s generalization across diverse deployment environments. To address this, we introduce the first generalized cross-domain few-shot (GCFS) task in 3D object detection, which focuses on adapting a source-pretrained model for high performance on both common and novel classes in a target domain with few-shot samples. Our solution integrates multi-modal fusion and contrastive-enhanced prototype learning within one framework, holistically overcoming challenges related to data scarcity and domain adaptation in the GCFS setting. The multi-modal fusion module utilizes 2D vision-language models to extract rich, open-set semantic knowledge. To address biases in point distributions across varying structural complexities, we particularly introduce a physically-aware box searching strategy that leverages laser imaging principles to generate high-quality 3D box proposals from 2D insights, enhancing object recall. To effectively capture domain-specific representations for each class from limited target data, we further propose a contrastive-enhanced prototype learning, which strengthens the model’s adaptability. We evaluate our approach with three GCFS benchmark settings, and extensive experiments demonstrate the effectiveness of our solution for GCFS tasks. The code will be publicly available.
1 Introduction
LiDAR-based 3D object detection [29, 23] has made considerable progress in recent years, largely driven by the availability of open-source datasets collected from various regions worldwide [10, 1, 36, 11]. These datasets have played a crucial role in advancing autonomous driving by providing foundational data for model training and evaluation. However, their coverage remains limited to a narrow range of object categories, often lacking the diversity needed to support generalization across various geographic and cultural contexts. Table 1 provides an overview of these datasets, highlighting the differences in object-detection focus. As models trained on these datasets are deployed in different regions or countries, they may encounter novel, region-specific objects that were not present in the original training data. Examples include electric scooters in China, tuk-tuks in Thailand, and indigenous wildlife in Africa.
In real-world applications of autonomous driving, new environments and novel object categories constantly emerge. Relying on large-scale datasets for each new environment to retrain or fine-tune a model from scratch is both time-consuming and resource-intensive, making it impractical for rapid deployment needs. Existing approaches on LiDAR-based 3D object detection, such as semi-supervised learning [41] and 3D open-vocabulary detection (OVD) [6, 51, 2], assume the availability of large amounts of unlabeled target-domain data, which may not always be feasible in deployment scenarios. Furthermore, while 3D-domain adaptation (DA)  [40, 48, 13] focuses on addressing domain shifts, it does not explicitly account for novel object categories that were unseen during training. Simply labeling novel objects as ”Others” is often insufficient in safety-critical scenarios where object-specific recognition is necessary for decision-making.
Figure 1: GCFS in 3D object detection aims to adapt the source-pretrained model for strong performance on common and novel classes in the target domain via limited target samples.
Table 1: Summary of common 3D Object Detection Datasets, where the most common detection categories are highlighted by blue.
Datasets Locations Classes Categories-of-interest
KITTI [10] Karlsruhe (Germany) $7$ Car, Pedestrian, Truck, Van, Person_sitting, Cyclist, Tram
NuScenes [1] Boston (USA), Singapore $23$ Car, Pedestrian, Truck, Bicycle, Barrier, Construction_vehicle, etc.
Waymo [36] 6 cities in USA $4$ Vehicle(including car, truck, and bus), Pedestrian, Cyclist, Sign
A2D2 [11] 50 cities in Germany $14$ Car, Pedestrian, Truck, Bicycle, Bus, UtilityVehicle, etc.
To bridge the gap from dataset-based training to real-world deployment, we define a new task, generalized cross-domain few-shot learning (GCFS), for LiDAR-based 3D object detection. As conceptualized in Fig. 1, the GCFS task explicitly addresses both cross-domain adaptation and few-shot learning for novel categories, offering a cost-effective solution for rapid deployment in diverse environments. Unlike existing 3D few-shot learning (FSL)  [53, 37, 18], or its extension, 3D generalized few-shot learning (GFSL) [19], which assume the same distribution between training and deployment environments, GCFS accommodates both domain discrepancies and novel object adaptation while maintaining robust performance on common categories.
Specifically, in the GCFS task, a 3D object detection model is initially trained with a source dataset that includes common object categories along with other possible source-specific classes. In the target deployment environment which may have domain gap from the source environment due to environmental factors, sensor configurations, and object appearances [48, 13, 17], we assume the presence of additional target-specific classes (i.e., novel classes) alongside the common classes. Given the practical feasibility and high cost of LiDAR data collection and annotation, we further assume that access to annotated data in the target environment is restricted to only a minimal amount (e.g., few-shot samples). The GCFS task, therefore, aims to enable the pre-trained model to adapt with minimal supervision in the target environment, ensuring strong performance on both common and target-specific novel categories. Though certain tasks in 2D object detection, like few-shot domain adaptation [9, 25] and generalized few-shot learning [7, 52], offer methodological insights into combining limited data adaptation with domain gap bridging, extending these 2D solutions effectively to the 3D domain remains challenging due to the higher-dimensional complexity and unique spatial characteristics of 3D data.
In this work, we introduce the first effective solution that integrates multi-modal fusion with contrastive-enhanced prototype learning to holistically overcome data scarcity and domain adaptation challenges in the GCFS task. Our multi-modal fusion module leverages vision-language models (VLMs) to extract open-set semantic information from point-cloud aligned images, enabling image-viewed 3D box proposals for improved object recall. However, accurate projection of 2D proposals into 3D space is challenged by point distribution biases that arise due to variations in object shapes and the inherent scanning nature of LiDAR sensors. To address this, we introduce a novel physically-aware 3D box searching strategy, which incorporates laser imaging principles to generate precise box proposals, improving localization across objects with diverse structural complexities. Though our image-guided fusion module significantly improves RPN’s proposal accuracy, challenges such as domain gaps, sparse point density, and limited data still hinder stable object feature learning through simple fine-tuning. To efficiently capture domain-specific representations for each object class in the target domain, we introduce contrastive-enhanced prototype learning, leveraging few-shot examples to encode object knowledge into flexible prototype embeddings. During inference, these prototypes act as anchors for feature refinement, adaptively aligning target-domain features with learned representations, thereby enhancing detection performance across domain shifts. In evaluation, we design three GCFS benchmark settings and conduct extensive experiments to illustrate the effectiveness of our solution. In sum, our contributions are as follows:
  • We formulate the first generalized cross-domain few-shot task in 3D object detection, comprehensively addressing the challenges in deploying dataset-pretrained models in real-world environments with diverse object categories.
  • We introduce the first GCFS solution for 3D object detection. It integrates multi-modal fusion with contrastive-enhanced prototype learning to overcome data scarcity and domain adaptation challenges holistically.
  • In our image-guided multi-modal fusion module, we propose a novel physical-aware box searching strategy for 2D-3D box proposal, boost the recall of objects.
  • We propose a novel contrastive-enhanced prototype learning strategy, enhancing model’s ability to swiftly adapt to the target domain with limited target data.
  • We establish three benchmark settings for evaluating GCFS tasks in 3D object detection, providing a comprehensive assessment framework for future exploration.
2 Related Works
2.1 LiDAR-based 3D Object Detection
LiDAR-based 3D object detection [29, 23] aims to locate and classify objects of interest from input point clouds. Its models have mainly developed through point-based, voxel-based, and point-voxed-based. Point-based models [27, 31, 34] incorporate raw points and the PointNet-based backbones for fine-grained representation at the point level, albeit with high computational demands. Voxel-based methods [46, 22, 5, 54] represent the point cloud within a structured voxel grid and utilize sparse convolution for feature extraction, offering a trade-off between computational efficiency and spatial resolution. Point-voxed-based methods [32, 33] combine both, achieving a balance between efficiency and representation resolution, but often coming with increased model complexity and computation.
2.2 Few-shot Learning in Object Detection
In object detection, FSL aims to enable models to detect objects-of-interest with limited/few-shot labeled samples. In 2D tasks, extensive studies [44] address the challenge of data scarcity by leveraging techniques like meta-learning [45, 30], transfer learning [39, 3], and data augmentation [42]. In 3D object detection, most works focus on indoor scenarios. Based on VoteNet [28], Zhao et al. [53] introduce a prototypical vote module for local features refinement and a prototypical head module for global feature enhancement. Tang et al. [37] design a VAE-based prototype learning on top of it, and Li et al. [18] further leverage contrastive learning to learn more refined prototypical representations. However, extending 3D indoor object detection methods to outdoor scenarios is challenging due to sparse point clouds at greater distances, dynamic objects, and varying lighting and weather conditions. Recently, Liuet al. [19] propose the first outdoor generalized FSL solution for novel class learning. Yet without dealing with domain gaps in cross-domain scenarios, it leads to limited performance on GCFS settings.
2.3 Domain Adaptation in 3D Object Detection
The study of domain adaptation in 3D object detection mainly focus on unsupervised or semi-supervised settings. Corral-Soto et al. [4] integrate CycleGAN into the detection model, leveraging target-domain data with a small fraction of labeled frames to adapt the model to the target domain. Yang et al. [48, 3] employ a hybrid quality-aware triplet memory to generate pseudo-labels for unlabeled target-domain data. Hegde et al. [13] propose a source-free unsupervised DA approach that utilizes class prototypes to suppress noisy pseudo-labels on target data. These methods rely on a large amount of target data for domain adaptation and cannot deal with novel classes in the target domain, making them inapplicable to the GCFS task.
2.4 Open-vocabulary 3D Object Detection
Recently, open-vocabulary object detection [43, 49, 12, 50, 15] is garnering significant attention. In 3D object detection, these methods usually leverage 2D VLMs to acquire novel open-set semantics and enable detection models in identifying novel objects without annotations [21, 51, 6] . For instance, Lu et al. [21] propose to utilize CLIP-based VLMs to connect open-set textual knowledge and point-cloud representations for novel object identification. Najibi et al. [24] auto-label point cloud sequences using a pretrained 2D VLM and enable open-set training via vision-language knowledge distillation. Zhang et al. [51] introduce 2D-3D co-modeling that estimates corresponding 3D boxes from 2D insights with temporal and spatial constraints. Etchegaray et al. [6] employ 2D VLMs to generate pseudo-labels, augmented with simulation data. Since these 3D-OVD methods rely on large volumes of target data (including novel objects), their performance on the GSFS task remains to be validated.
3 Preliminaries: Generalized Cross-domain Few-Shot Learning
Figure 2: Proposed GCFS Framework. We first pretrain a detection model with source data. During model finetuning using target few-shot data, pairs of image and point cloud are processed by GDino+SAM and 3D backbone for image masks and 3D features. Insights from 2D analysis contribute in 1) enriching 3D features with 2D context and 2) proposing high-quality proposals via novel 2D-to-3D box search. Proposal features after RoI pooling are refined by learnable prototypes with attention mechanism, and then passed to final prediction.
Generalized Cross-domain Few-shot 3D object detection is a novel task designed to address the challenges of deploying 3D object detection models in real-world scenarios, where models trained on a source dataset must adapt to a new, target environment with limited labeled data across diverse categories. It is defined by two primary constraints. First, it assumes that only a limited number of labeled samples (i.e., few-shot samples) are available for all classes in the target environment, including both common classes (those shared with the source domain) and novel, target-specific classes. This reflects the practical limitations in collecting extensive data for every class in a new environment. Second, the GCFS setting accounts for a domain gap between the source and target environments, which may differ significantly in terms of object appearance, environmental conditions, and sensor setups. This gap presents an additional challenge for generalizing the model’s performance. Compared to FSL, DA, and OVD tasks, GCFS requires the solution to holistically address challenges of data scarcity and domain gap in real-world deployment of 3D object detection.
To formulate GCFS of 3D object detection, we distinguish LiDAR data from the source dataset and target environment (dataset) with superscript $s$ and $t$, respectively. In the source dataset used to pre-train a detection model $M_{pretrained}$, we assume sufficient amount of annotated data $D^{s}=\{\textbf{B}^{s}_{i},\textbf{C}^{s}_{i},\textbf{P}^{s}_{i}\mid\textbf{P}_{i}^{s}\in\mathbb{R}^{N\times 3}\}$ is available for the $\textbf{C}^{s}$ categories of objects-of-interest, where $\textbf{B}^{s}_{i}=\{\textbf{b}^{s}\mid\textbf{b}^{s}=[x,y,z,h,w,l,\theta]\}_{l=1}^{L_{i}}$ are the box annotations in point cloud $\textbf{P}_{i}^{s}$. In terms of target dataset $D^{t}=\{\textbf{B}^{t}_{i},\textbf{C}^{t}_{i},\textbf{P}^{t}_{i}\}$, only limited amount of data (e.g. few-shots) of the $\textbf{C}^{t}$ categories of object-of-interest is available. Here, we assume some categories in $\textbf{C}^{t}$ same as those in $\textbf{C}^{s}$, so certain knowledge in the pretrained model is valuable to the target task. In mathematics, these common classes are defined by $C_{com}=\textbf{C}^{t}\cap\textbf{C}^{s}\neq\emptyset$. We use $C_{nov}^{s}=\textbf{C}^{s}\setminus C_{com}$ and $C_{nov}^{t}=\textbf{C}^{t}\setminus C_{com}$ to denote the domain-specific classes, or novel classes. That is, objects in $C_{nov}^{t}$ are unseen in the source dataset. The target of the GCFS task is to obtain a strong detection model $M_{finetuned}$ through refining the model $M_{pretrained}$ with the $K$-shot examples in $D^{t}$.
4 Methodology
Fig. 2 illustrates the systematic diagram of our holistic solution for addressing the GCFS problem in LiDAR-based 3D object detection. It integrates multi-modal fusion (4.1) with contrastive-enhanced prototype learning (4.2) to overcome data scarcity and domain adaptation challenges. Briefly, our solution begins by pre-training a standard 3D object detection model on source data. In the target domain finetuning, a multi-modal fusion module leverages VLMs to extract rich, open-set semantic features from images. We specifically design a physically-aware box searching module to supplement the 3D region proposal network (RPN). Furthermore, to efficiently capture domain-specific representations for each class, we introduce contrastive-enhanced prototype learning supported by the few-shot examples. During inference, these prototypes serve as anchors for feature refinement, allowing the model to adaptively align target features with learned representations, thereby strengthening detection robustness and accuracy across domain gaps. This combination of multi-modal fusion and prototype-based feature refinement enables the model to generalize effectively to the target domain with minimal labeled data, addressing the unique requirements of the GCFS setting for real-world 3D object detection applications.
4.1 Image-guided Multi-modal Fusion
Trained on large-scale datasets, 2D Region VLMs [16, 14, 20] offer strong generalization capabilities to understand 2D visual information at a fine-grained, region-based level. In our GCFS task, we leverage point-cloud-aligned images and 2D VLMs, specifically, SAM [14] and Ground DINO (GDino) [20], to supplement the standard 3D detection. It not only enhances object detection (especially those in the novel classes) under the few-shot constraint, but also helps bridge the domain gap between source and target environments. In specifical, our image-guided multimodal fusion enhances the RPN’s object recall in twofolds. First, numerical features from the image analysis branch are integrated into the point-cloud representation, enriching it with 2D contextual information. Second, the 2D bounding boxes and masks are re-projected into 3D space to align with the point cloud data through our innovative physical-aware box searching. This re-projection addresses biases in point distributions across different object shapes that is often overlooked by existing box searching methods [6, 51]. By leveraging principles of laser imaging, our physical-aware 3D box searching strategy accurately generates 3D box proposals for objects of varying structural complexities.
Image-point feature fusion. In the point cloud branch, point cloud P is passed through the backbone encoder for point feature $\textbf{F}^{3D}\in\mathbb{R}^{N_{voxel}\times C}$. Simultaneously in the image branch, given the input image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, we take object categories (including $C_{com}$ and $C_{nov}^{t}$) as text prompts to trigger GDino for detecting potential objects of interest with 2D boxes $\textbf{B}^{2D}\in\mathbb{R}^{N_{obj}\times 4}$. After the Non-maximum suppression (NMS) and score filtering, SAM takes $\textbf{B}^{2D}$ as box prompts to generate the object mask $\textbf{M}^{2D}\in\mathbb{R}^{H\times W\times|C^{t}|}$. Then, we utilize 2D-3D projection transformation on the coordinates of voxels $\textbf{P}^{voxel}$ to find corresponding 3D voxels of $\textbf{M}^{2D}$, obtaining voxel-wise feature $\textbf{F}^{voxel}\in\mathbb{R}^{N_{voxel}\times|C^{t}|}$,
$\textbf{F}^{voxel}=f_{proj}(\textbf{M}^{2D},\textbf{F}^{3D},\textbf{P}^{voxel}),$ (1)
where $f_{proj}$ denotes the re-projection function, defined by the intrinsic of the camera and extrinsic between the camera and LiDAR. To enrich the 3D features with 2D context information, we utilize a simple MLP to align channels of $\textbf{F}^{voxel}$ and $\textbf{F}^{3D}$ for a comprehensive feature $\textbf{F}^{fused}$,
$\textbf{F}^{fused}=\textbf{F}^{3D}+\textit{MLP}(\textbf{F}^{voxel})$ (2)
This feature fusion improve the following RPN for a better recall on objects of interest.
Physical-aware box searching. To find fine-grained box location from the 2D masks, we first take raw points as input to identify points within image masks,
$^{pts}=f_{proj}(\textbf{M}^{2D},\textbf{P}),$ P ) , (3)
where $\textbf{P}^{pts}$ is the point set corresponding to $\textbf{B}^{2D}$. We denote the point cloud of the $i^{th}$ object in $\textbf{B}^{2D}$ as $\textbf{P}_{i}\in\textbf{P}^{pts}$, and use the mean and 2×2$\times$2 ×-standard-deviation of all these point coordinates as the center and the boundary of the valid range to wipe out possible background points. For each class $c\in\textbf{C}^{t}$, we pre-define an anchor box $[h^{c},w^{c},l^{c}]$ according to mean object sizes of target few-shot data. The goal of box searching is to find the optimal location $[x,y,z]$ and heading angle $\theta$ for each $\textbf{P}_{i}$. Specifically, the variables $[x,y,z,\theta]$ shapes a reverse transformation T from $\textbf{P}_{i}$ to local coordinates $\textbf{P}^{local}_{i}$, $\textbf{P}^{local}_{i}=\textbf{T}\textbf{P}_{i}$, where
$\textbf{T}=\begin{bmatrix}\cos\theta&-\sin\theta&0&t_{y}\sin\theta-t_{x}\cos\theta\\ \\ \sin\theta&\cos\theta&0&-t_{x}\sin\theta-t_{y}\cos\theta\\ 0&0&1&-z\\ \end{bmatrix}.$ (4)
To find the right set of $[x,y,z,\theta]$, we design an outside distance loss $L_{OD}$ to ensure $\textbf{P}^{local}_{obj}$ are inside the box,
$L_{OD}=\sum_{c=1}^{C_{cls}}\sum_{\textbf{P}^{local}_{obj}\in c}min(|\textbf{P}^{local}_{i}|-\textbf{BD}^{c},0).$ (5)
Here, $\textbf{BD}^{c}=[h^{c}/2,w^{c}/2,l^{c}/2]$ denotes the boundary of local points from class $c$.
Figure 3: Physical-aware box searching, where red boxes are GT boxes and blue ones are searched boxes. Regarding “Cyclist” (left) and “Car” (right), searched boxes have biased angles and centers, respectively, which are corrected by $L_{BVC}$ and $L_{FVD}$.
Furthermore, we notice that due to central, unidirectional scanning, LiDAR-based object points present significant differences in point distribution regarding different structural complexities. For simple structural objects with flat surfaces, mostly including vehicles (e.g., cars, trucks, and buses), most points are on the smooth surfaces, which are front-viewed by LiDAR. For complex structural objects with irregular surfaces (e.g., pedestrians, bicycles, and sitting persons), points are more to shape the whole objects in the bird’s view. Motivated by this observation, we categorize general objects into two types by their structural complexity: simple structural (SS) objects and complex structural (CS) objects. For SS objects, we design the front-viewed distance (FVD) loss to make points closer to the front-viewed boundaries of boxes,
$L_{FVD}=\sum_{\begin{subarray}{c}c=1\\ c\in SS\end{subarray}}^{C_{cls}}\sum_{\textbf{P}^{local}_{i}\in c}(|\textbf{P}^{local}_{i}|-\textbf{FB}^{c}),$ (6)
where $\textbf{FB}^{c}$ is LiDAR-front-viewed box boundaries determined by the box heading angle $\theta$ and box location $[x,y,z]$ (refer to Supplementary for details). For CS objects, we design the bird-viewed center (BVC) loss to align the centers of points and boxes.
$L_{BVC}=\sum_{\begin{subarray}{c}c=1\\ c\in CS\end{subarray}}^{C_{cls}}\sum_{\textbf{P}^{local}_{i}\in c}|f_{2D}(\textbf{P}^{local}_{i})|,$ (7)
where $f_{2D}$ is the function to obtain $[x,y]$ of $\textbf{P}^{local}_{obj}$. Applying $L_{BVC}$ and $L_{FVD}$ facilitates finding correct centers and heading angles for boxes, as shown in Figure 3.
In summary, the loss of box searching for $[x,y,z,\theta]$ is:
$L_{box}=L_{OD}+\lambda_{1}L_{FVD}+\lambda_{2}L_{BVC}$ (8)
Since the computational load of box searching is low (due to the limited number of object points) and does not involve in model parameter update, we use the Quasi-Newton ”BFGS” function of the Scipy tool to quickly calculate the set of location $[x,y,z]$ and heading angle $\theta$ for each $\textbf{P}^{local}_{i}$.
4.2 Contrastive-Enhanced Learnable Prototype and Feature Refinement
Our image-guided fusion significantly improves RPN proposal accuracy. Yet, the domain gap, sparse point density, and data limitation still prevent the model from learning reliable object features through simple model fine-tuning. To tackle this, we design a novel contrastive-enhanced prototype learning strategy from the limited samples. Refining proposal features with such prototypes enhances the model’s ability to detect objects-of-interest. Different from [18] that uses contrastive learning to enhances static prototypes, our approach leverages few-shot-driven contrastive learning on learnable prototypes, making our learnable prototypes more discriminative and comprehensive.
Figure 4: Few-shot feature extraction and CL-enhanced prototype learning. (In few-shot feature extraction, 2D and 3D ground-truth labels replace GDINO and RPN outputs to obtain object features.)
Contrastive-enhanced prototype learning. We build a learnable targe-specific feature bank $\textbf{F}^{pro}\in\mathbb{R}^{|C^{t}|\times C}$ for all object classes. These prototypes are optimized together with the model finetuning update. To accelerate their converge under limited data constraint, we introduce a contrastive loss directly on these learnable prototypes. To this end, as shown in Figure 4, we pool few-shot object features $\textbf{F}^{fs}$ according to their ground-truth annotation as the contrastive anchors. Then for each class $c\in C^{t}$, we construct positive pairs with the corresponding prototype $\textbf{F}^{pro}_{c}$ and its anchor $\textbf{F}^{fs}_{c}$. The rest prototypes in the feature bank, denoted by $\textbf{F}^{pro}_{s}$ for $c\neq s$, are negative samples of the anchor.
$L_{CL}=-\sum_{c\in C_{cls}}\log\frac{\exp(\text{Sim}(\textbf{F}^{fs}_{c},\textbf{F}^{pro}_{c}))}{\sum_{s\in C_{cls}}\exp(\text{Sim}(\textbf{F}^{fs}_{c},\textbf{F}^{pro}_{s}))},$ (9)
where $\text{Sim}(\cdot)$ calculates the cosine similarity between two numerical features in the InfoNCE contrastive loss [26]. Since the anchors are directly obtained from examples in the target domain, our contrastive-enhanced features helps bridge the domain gap between source and target environments and speed up $\textbf{F}^{pro}$ acquiring semantic essences of various classes under limited training data.
Feature refinement by prototypes. After obtaining the $\textbf{F}^{pro}$ along with model finetuning process, we freeze the prototypes and use them to refine the proposal features $\textbf{F}^{prp}$ from the query data. In the multi-head cross-attention, we take $\textbf{F}^{pro}$ to form the key and value and $\textbf{F}^{prp}$ as the query.
$\hat{\textbf{F}}^{prp}=\text{Softmax}(\frac{{\textbf{F}}^{prp}\textbf{W}_{Q}({\textbf{F}}^{pro}\textbf{W}_{K})^{T}}{\sqrt{d}}){\textbf{F}}^{pro}\textbf{W}_{V},$ (10)
where $d$ is the dimension of $\textbf{F}^{prp}$, and $\textbf{W}_{Q},\textbf{W}_{K},\textbf{W}_{V}$ are trainable transformations of query, key, and value. Finally,
$\tilde{\textbf{F}}^{prp}=\hat{\textbf{F}}^{prp}+\textbf{F}^{prp}.$ (11)
is passed to the object detection head for object detection.
4.3 Model Optimization and Inference
The 3D detection model parameter update and our prototype learning are collectively conducted together. The overall loss $L$ to optimize the whole system is:
$L=L_{rpn}+L_{det}+{\lambda}L_{CL}$ (12)
where $L_{rpn}$ and $L_{det}$ are the standard detection loss of RPN and detection head, and $\lambda$ is a weight hyper-parameter. To further enable the model to efficiently adapt to a new domain under the limited data setting, we adopt a MAML-based [8] training scheme. Briefly, during meta-training, we leverage the source data to set up the $K$-shot meta-task.This meta-training facilitates to find a set of model parameters and $\textbf{F}^{pro}$ for the quick model adaptation in the unseen target domain. Please refer to the supplementary for more details.
During model deployment, the query point cloud and corresponding 2D image undergo the proposed image-guided fusion to enhance voxel representation and proposal boxes. After ROI pooling, the extracted features are further refined with class prototypes to improve discrimination.
5 Experimentation
Table 2: Performance in mAP(%) for NuScenes $\rightarrow 5$shot-KITTI, Waymo $\rightarrow 5$shot-KITTI, and KITTI $\rightarrow 5$shot-A2D2. The bold values represent the best performance. Subscript values in parentheses are standard deviations. Please refer to the supplementary for specific detection performance across various categories.
Methods Venues NuScenes→FS-KITTI Waymo→FS-KITTI KITTI→FS-A2D2
common novel overall common novel overall common novel overall
VoxelRCNN Target-FT - 12.77(1.85) 5.48(1.16) 8.61(1.45) 23.06(1.64) 12.47(1.92) 17.01(1.8) 5.09(1.1) 0.7(0.17) 2.9(0.64)
Proto-Vote [53] NIPS’22 7.56(2.39) 5.74(1.54) 6.52(1.9) 17.36(2.72) 12.08(1.78) 14.34(2.19) 3.61(0.81) 1.86(0.48) 2.74(0.65)
PVAE-Vote [37] NIPS’24 8.01(2.77) 6.38(2.22) 7.08(2.46) 18.19(2.94) 12.79(2.2) 15.1(2.52) 3.43(0.85) 1.97(0.54) 2.7(0.7)
CP-Vote [18] PRCV’24 10.69(2.28) 7.84(1.87) 9.06(2.04) 17.66(2.4) 12.17(1.86) 14.52(2.09) 4.28(0.92) 2.72(0.88) 3.5(0.9)
GFS-Det [19] arXiv’23 12.83(2.36) 1.18(0.38) 6.17(1.23) 22.74(2.78) 1.26(0.38) 10.47(1.41) 4.39(0.57) 0.22(0.05) 2.3(0.31)
Ours - 15.99(1.62) 11.72(1.4) 13.55(1.5) 25.4(2.02) 17.75(1.66) 21.03(1.82) 7.78(0.68) 5.22(0.61) 6.5(0.64)
PV-RCNN++ Target-FT - 15.28(2.03) 6.39(1.71) 10.2(1.85) 24.35(1.4) 13.79(1.73) 18.31(1.59) 3.64(1) 0.26(0.09) 1.95(0.55)
Proto-Vote [53] NIPS’22 6.97(1.43) 7.05(2.02) 7.02(1.76) 10.02(2.16) 12.27(1.49) 11.31(1.78) 3.92(0.81) 1.48(0.5) 2.7(0.65)
PVAE-Vote [37] NIPS’24 7.43(1.72) 7.53(2.16) 7.49(1.97) 11.36(2.4) 12.84(1.93) 12.21(2.13) 3.97(0.96) 1.34(0.54) 2.66(0.75)
CP-Vote [18] PRCV’24 8.73(1.97) 7.17(1.8) 7.84(1.87) 14.18(2.79) 13.32(2.74) 13.69(2.76) 4.2(1.1) 2.14(0.63) 3.17(0.87)
GFS-Det [19] arXiv’23 17.37(2.08) 1.16(0.14) 8.1(0.97) 24.79(1.79) 1.33(0.23) 9.67(0.9) 3.46(0.56) 0.14(0.04) 1.8(0.3)
Ours - 18.06(1.46) 11.11(1.03) 14.09(1.21) 27.35(1.55) 16.02(1.93) 20.88(1.77) 9.75(1.09) 5.76(0.8) 7.76(0.94)
5.1 Experimental Settings [^1] More details on experimental settings (e.g. benchmark setup, implementation details, and comparison methods) are provided in our supplementary.
[…]
Implementation details. Voxel-absed VoxelRCNN [5] and point-voxel-based PVRCNN++ [33] are taken as the base detection models in experiments and our solution is implemented based on OpenPCDet [38] and RoboFusion [35]. Data augmentations during model pre-training include random world flipping, random world scaling, and random world rotation. In model fine-tuning, we additionally involve ground-truth object sampling to ensure the existence of all classes in each iteration. For box searching, we define CS classes [Pedestrian, Person_sitting, Cyclist, Bicycle, Utility_vehicle] and SS classes [Car, Truck, Van, Tram, Bus]. The “Adam-Onecycle” optimizer [38] with a learning rate of 0.01 is adopted. Epochs of pre-training are 30 for NuScenes and Waymo, and 80 for KITTI, the epoch of fine-tuning is 100 for both tasks. All experiments are conducted on a machine with dual-GeForce RTX-3090.
Compared Methods. As no prior work has specifically tackled the GCFS task for outdoor 3D object detection, we use a simple fine-tuning approach on target-domain data (denoted as Target-FT) as the baseline. To benchmark our method, we compare against SOTA 3D-FSL methods, including Proto-Vote [53], PVAE-Vote [37], and CP-Vote [18], as well as the 3D-GFSL method GFS-Det [19]. Note that current outdoor OVD methods (i.e., Unsup3D [24], FnP [6], and OpenSight [51]) and 3D-DA methods (i.e., SN [40] and ST3D++ [48, 47]) are not directly applicable to our GCFS benchmark, as they rely on extensive data, additional temporal information, or unsupervised learning without annotation-based supervision. To further assess the generalizability and potential of our approach, we extend our ablation study to a more complex unsupervised few-shot learning setting, where these 3D-OVD and 3D-DA methods can be evaluated under conditions more aligned with their original assumptions.
5.2 Experimental Results on GCFS Benchmark
For the $5$-shot GCFS tasks, our proposed method outperforms all compared approaches, as shown in Table 2. It achieves higher mAP for both common and novel categories [^3] The specific detection performance across various categories on the three benchmark settings is reported in the supplementary., demonstrating strong knowledge transferability from the source domain while effectively adapting to novel classes with limited supervision. Furthermore, our method exhibits robust adaptability across both high-density (64-beam) KITTI and low-density (16-beam) A2D2, ensuring robust few-shot detection even in challenging low-density LiDAR scenarios [17]. In contrast, existing 3D-FSL and 3D-GFSL methods struggle in GCFS tasks, primarily due to their limited ability to bridge domain gaps. For the 3D-FSL methods, their reliance on dense, close-range point clouds further degrades performance when detecting sparse, distant objects under varying lighting and weather conditions. GFS-Det particularly struggles with novel categories due to its simplistic incremental learning strategy, which prevents effective knowledge transfer from common classes to novel ones. In addition, we observe that all methods exhibit lower performance in the KITTI (64-beam) → FS-A2D2 (16-beam) setting. The significant reduction in LiDAR resolution introduces a larger domain gap, making domain adaptation more challenging compared to other settings [17].
Limitation on hard classes. Although our method outperforms other methods (see supplementary for details), it shows limited improvement on certain hard classes, (i.e., “Person_sitting”) due to the diverse and misleading structures of these samples, making feature learning more challenging. In the future, we plan to enhance our method to better handle these hard classes.
5.3 Ablation Studies
We conduct ablation experiments mainly on NuScenes $\rightarrow 5$shot-KITTI with VoxelRCNN as detection model, to further analyze our method (see the supplementary for details).
Table 3: Component ablations in mAP(%). (Image-Fusion is our proposed image-guided multi-modal fusion and CL-Proto is our proposed contrastive-learning-enhanced prototype learning.)
Target-
FT
Image-
Fusion
CL-
Proto
Common Novel Overall
(a) 12.77 5.48 8.61
(b) 14.80 8.10 10.97
(c) 14.69 11.17 12.68
(d) 15.99 11.72 13.55
Table 4: Performance in mAP(%) of adopting prototype learning with or without contrastive learning (CL).
Prototype Common Novel
w/o CL 15.23 10.33
w/ CL 15.99 11.72
Table 5: Performance in mAP(%) with box searching by $L_{OD}$ only or $L_{box}$ (w/ $L_{OD}$, $L_{FVD}$, and $L_{BVC}$).
Box Search CS SS
$L_{OD}$ 4.65 12.56
$L_{box}$ 6.74 13.58
Component Ablation. Table 3 indicates that the proposed contrastive-enhanced prototype learning significantly enhances performance across overall classes with a notable boost for common classes, as (a)$\rightarrow$(b), illustrating its swift adaptation to limited samples in target domain. When applying our image-guided multi-modal fusion, the detection model shows marked improvement, especially on novel classes, as (a)$\rightarrow$(c), showing its boost on object recall. By combining both, our GCFS method achieves the highest performance, showing their complementary to each other.
We also conduct ablations on our proposed prototype learning and box searching components. Table 5 shows that the contrastive loss boosts model performance, indicating its ability to swiftly adapt prototypes to few-shot data in target domain. In Table 5, integrating $L_{OD}$ with $L_{FVD}$ and $L_{BVC}$ yields improvements on both CS and SS objects, showing $L_{FVD}$ and $L_{BVC}$ enhancing recall rates for objects with diverse structural complexities, thereby further optimizing model effectiveness.
Ablation on different shots. From Table 6, we observe, as K increases, our method demonstrates strong scalability to multi-shot settings, effectively improving both common and novel class detection. At K=40, our overall performance (26.05%) closely approaches full-shot (28.21%), bridging the gap between few-shot and full supervision. Notably, our novel-class performance surpasses full-shot training, due to data imbalance in full-shot learning and our image-guided approach enhancing novel object discovery. These results confirm that our method scales well with more shots, achieving strong generalization even with limited data.
Table 6: Performances in mAP(%) with different $K$. (Full-shot denotes the training on the complete KITTI train set under pre-training settings.)
K Common Novel Overall
1 7.27 0.57 3.44
3 12.27 7.76 9.70
5 15.99 11.72 13.55
10 23.56 12.21 17.08
20 27.59 17.49 21.82
40 32.05 21.55 26.05
Full-shot 41.34 18.35 28.21
Table 7: Comparison in mAP(%) with OV and DA methods in the unsupervised few-shot setting. (Note that Unsup3D [24] and OpenSight [51] require sequential point clouds as inputs, which doesn’t apply to the discrete few-shot target data in GCFS.)
Methods Venue Common Novel Overall
DA SN [40] CVPR’20 12.09 - -
ST3D++ [48, 47] PAMI’22 21.00 - -
OVD FnP [6] ECCV’24 10.59 2.66 6.06
Ours-OVD - 22.25 8.26 14.26
Unsupervised few-shot ablation with OV and DA methods. We establish an unsupervised few-shot setting with no annotations for all classes. Our approach is benchmarked against SOTA 3D-OVD solution, FnP [6], and well-established 3D-DA methods, SN [40] and ST3D++ [48, 47]. To create an OVD version of our method, we incorporate a physical-aware box searcher to generate high-quality pseudo-labels for target-specific training. Table 7 shows that, compared to OVD and DA methods, our OVD method achieves the highest mAPs, showing strong domain gap bridging capability and high learning efficiency from unlabeled samples. Please refer to the supplementary for implementation and result details.
6 Conclusion
This paper proposed the generalized cross-domain few-shot task in 3D object detection, aiming to adapt a source-pretrained model for strong performance on common and novel classes in the target domain with few-shot samples. We introduced the first GCFS solution, featuring a multi-modal fusion module for feature fusion and box proposal searching and contrastive-enhanced prototype learning for quick feature adaptation. In the established GCFS benchmarking settings, extensive experiments validated the effectiveness of our solution.
Supplementary Material
In this supplementary material, we provide additional details on our proposed image-guided multi-modal fusion module and the optimization-based meta-learning scheme in Section 7. In addition, we provide further details on the benchmark settings and method implementations, along with a comprehensive presentation of the experimental results in Section 8.
7 Details on Methodology
7.1 Image-guided Multi-modal Fusion
Front-viewed boundary calculation. Regarding the FVD loss, we utilize the LiDAR’s view angle to obtain the front-reviewed boundary FB for each class. Specifically, for given box $[x,y,z,\theta]$, we first get the yaw angle by $\alpha=\arctan(\frac{x}{y})$. Then, we obtain the view angle $\phi=\alpha-\theta$, which indicates the LiDAR scanning direction w.r.t. the search box. Regarding $\phi$, we define FB via the prior box length $l$ and width $w$:
$\text{{FB}}=\left[\frac{l}{2}S_{l},\frac{w}{2}S_{w}\right],\\ \text{s.t. }[S_{l},S_{w}]=\begin{cases}\left[-1,-1\right],&0<\phi\leq\frac{\pi}{2}\\ \left[1,-1\right],&\frac{\pi}{2}<\phi\leq\pi\\ \left[1,1\right],&\pi<\phi<\frac{3\pi}{2}\\ \left[-1,1\right],&\frac{3\pi}{2}<\phi\leq 2\pi\end{cases}$ (13)
which indicates the faces of the box front-viewed by scanning LiDARs.
7.2 Optimization-based Meta-learning
In meta-training, we utilize the sufficient data of common classes $C_{com}$ and other classes $C_{bas}^{s}$ to simulate the target few-shot fine-tuning on $C_{com}$ and $C_{nov}^{t}$. Specifically, we first randomly sample $N_{nov}$ classes from $C_{bas}^{s}$, where $N_{nov}$ is the class number of $C_{nov}^{t}$. Then, as shown in Figure 5, in the inner loop, we set up a $K$-shot cross-domain detection task, covering common classes and sampled $N_{nov}$ classes. For the meta-task, we utilize data augmentations on support data to simulate domain gaps regarding source-trained prototypes transferred to target data. Liked in MAML [8], we design the outer loop, where we run one inner loop to get the one-step updated parameter, run another inner loop to get the two-step gradient, and use the two-step gradient to update the original model parameters. As in [8], this inner-outer-loop meta-learning will find a set of parameters and prototypes $\textbf{F}_{proto}$ with the quick adaptation to a new few-shot learning task (covering common classes and novel classes) in a new domain via limited target data.
Figure 5: Meta-learning scheme simulating few-shot learning with domain gaps.
Table 8: Performance comparison in mAP(%) on VoxelRCNN detector in the NuScenes $\rightarrow 5$shot-KITTI GCFS task. The bold values represent the best performance. Subscript values in parentheses are standard deviations. The ped is for Pedestrian, trk for Truck, ps for Person_sitting, cyc for Cyclist, and trm for Tram.
Methods Venues car ped trk common van ps cyc trm novel overall
Target-FT - 30.08(2.75) 7.08(1.86) 1.17(0.94) 12.77(1.85) 14.24(1.61) 0.87(0.3) 4.59(2.01) 2.24(0.71) 5.48(1.16) 8.61(1.45)
Proto-Vote NIPS’22 17.93(5.88) 4.49(1.07) 0.26(0.22) 7.56(2.39) 11.36(3.19) 0.02(0.02) 9.86(2.14) 1.7(0.8) 5.74(1.54) 6.52(1.9)
PVAE-Vote NIPS’24 18.76(6.17) 4.06(1.64) 1.23(0.51) 8.01(2.77) 12.07(3.67) 0.05(0.04) 11.22(3.17) 2.17(2) 6.38(2.22) 7.08(2.46)
CP-Vote PRCV’24 24.7(4.78) 6.34(1.51) 1.02(0.54) 10.69(2.28) 11.2(3.16) 0.57(0.21) 15.5(2.52) 4.08(1.59) 7.84(1.87) 9.06(2.04)
GFS-Det arXiv’23 17.1(3.27) 19.86(3.67) 1.54(0.13) 12.83(2.36) 2.42(0.55) 0.05(0.02) 2.15(0.92) 0.1(0.02) 1.18(0.38) 6.17(1.23)
Ours - 37.71(2.41) 7.16(1.89) 3.11(0.56) 15.99(1.62) 22.29(2.43) 1.54(0.28) 18.26(1.56) 4.79(1.33) 11.72(1.4) 13.55(1.5)
Table 9: Performance comparison in mAP(%) on PVRCNN++ detector in the NuScenes $\rightarrow 5$shot-KITTI GCFS task. The bold values represent the best performance. Subscript values in parentheses are standard deviations. The ped is for Pedestrian, trk for Truck, ps for Person_sitting, cyc for Cyclist, and trm for Tram.
Methods Venues car ped trk common van ps cyc trm novel overall
Target-FT - 38.79(4.5) 6.13(0.8) 0.92(0.79) 15.28(2.03) 11.3(2) 2.08(1.29) 11.16(2.7) 1.05(0.86) 6.39(1.71) 10.2(1.85)
Proto-Vote NIPS’22 15.83(3.21) 3.09(0.25) 2(0.82) 6.97(1.43) 13.26(1.9) 0.61(0.4) 12.55(5.09) 1.78(0.66) 7.05(2.02) 7.02(1.76)
PVAE-Vote NIPS’24 17.1(3.22) 3.17(0.99) 2.02(0.95) 7.43(1.72) 14.12(2.18) 0.28(0.16) 13.67(5.48) 2.05(0.81) 7.53(2.16) 7.49(1.97)
CP-Vote PRCV’24 22.92(4.26) 2.74(1.26) 0.53(0.41) 8.73(1.97) 12.55(2.37) 0.55(0.31) 13.31(3.96) 2.29(0.56) 7.17(1.8) 7.84(1.87)
GFS-Det arXiv’23 21.04(2.2) 30.16(3.77) 0.9(0.28) 17.37(2.08) 3.41(0.31) 0.05(0.01) 1.03(0.21) 0.13(0.02) 1.16(0.14) 8.1(0.97)
Ours - 44.92(2.61) 6.13(0.67) 3.14(1.09) 18.06(1.46) 21.16(1.35) 0.86(0.11) 18.05(2.03) 4.36(0.61) 11.11(1.03) 14.09(1.21)
Table 10: Performance comparison in mAP(%) on VoxelRCNN detector in the Waymo $\rightarrow 5$shot-KITTI GCFS task. The bold values represent the best performance. Subscript values in parentheses are standard deviations. The ped is for Pedestrian, trk for Truck, ps for Person_sitting, cyc for Cyclist, and trm for Tram.
Methods Venues car ped trk common van ps cyc trm novel overall
Target-FT - 55.86(2.93) 11.79(1.08) 1.52(0.89) 23.06(1.64) 21.03(3.49) 2.71(0.83) 25.54(3.04) 0.59(0.32) 12.47(1.92) 17.01(1.8)
Proto-Vote NIPS’22 39(2.07) 9.45(4.09) 3.64(2.01) 17.36(2.72) 20.24(2.77) 1.26(0.52) 23.79(2.33) 3.02(1.51) 12.08(1.78) 14.34(2.19)
PVAE-Vote NIPS’24 41.99(2.91) 9.55(4.12) 3.01(1.79) 18.19(2.94) 22.51(4.17) 1.01(0.48) 24.22(2.51) 3.41(1.64) 12.79(2.2) 15.1(2.52)
CP-Vote PRCV’24 41.35(2.5) 7.67(1.78) 3.95(2.92) 17.66(2.4) 20.42(2.31) 1.41(0.45) 22.63(3.09) 4.22(1.59) 12.17(1.86) 14.52(2.09)
GFS-Det arXiv’23 21.96(2.99) 42.83(4.37) 3.44(0.98) 22.74(2.78) 2.13(0.79) 1.04(0.3) 1.77(0.39) 0.11(0.04) 1.26(0.38) 10.47(1.41)
Ours - 57.36(1.78) 15.12(3.04) 3.73(1.26) 25.4(2.02) 28.7(1.69) 1.47(0.18) 30.48(2.32) 10.36(2.46) 17.75(1.66) 21.03(1.82)
Table 11: Performance comparison in mAP(%) on PVRCNN++ detector in the Waymo $\rightarrow 5$shot-KITTI GCFS task. The bold values represent the best performance. Subscript values in parentheses are standard deviations. The ped is for Pedestrian, trk for Truck, ps for Person_sitting, cyc for Cyclist, and trm for Tram.
Methods Venues car ped trk common van ps cyc trm novel overall
Target-FT - 59.18(2.88) 12.12(0.76) 1.74(0.55) 24.35(1.4) 24.94(2.73) 2.25(0.75) 26.88(2.57) 1.08(0.86) 13.79(1.73) 18.31(1.59)
Proto-Vote NIPS’22 22.36(3.93) 4.24(0.71) 3.45(1.84) 10.02(2.16) 22.12(2.03) 0.83(0.43) 23.76(2.99) 2.39(0.52) 12.27(1.49) 11.31(1.78)
PVAE-Vote NIPS’24 25.42(4.51) 5(0.79) 3.67(1.89) 11.36(2.4) 23.05(2.67) 1.26(0.72) 23.92(3.52) 3.13(0.82) 12.84(1.93) 12.21(2.13)
CP-Vote PRCV’24 38.09(6.66) 3.15(1.04) 1.31(0.68) 14.18(2.79) 22.3(3.93) 1.2(0.29) 25.94(6.01) 3.83(0.71) 13.32(2.74) 13.69(2.76)
GFS-Det arXiv’23 26.09(1.58) 43.02(2.55) 5.25(1.25) 24.79(1.79) 1.93(0.33) 1.23(0.3) 1.99(0.23) 0.16(0.05) 1.33(0.23) 11.38(0.9)
Ours - 59.2(1.8) 15.17(1.56) 7.68(1.3) 27.35(1.55) 28.06(3.02) 1.7(0.44) 27.01(2.79) 7.3(1.46) 16.02(1.93) 20.88(1.77)
Table 12: Performance comparison in mAP(%) on VoxelRCNN detector in the KITTI $\rightarrow 5$shot-A2D2 GCFS task. The bold values represent the best performance. Subscript values in parentheses are standard deviations. The trk for Truck, bcy is for Bicycle, and uvc for Utility_vehicle.
Methods Venues car ped trk common bcy uvc bus novel overall
Target-FT - 2.82(0.43) 3.58(1.06) 8.89(1.8) 5.09(1.1) 0.25(0.08) 0.04(0.03) 1.83(0.4) 0.7(0.17) 2.9(0.64)
Proto-Vote NIPS’22 1.89(0.27) 3.38(0.69) 5.58(1.49) 3.61(0.81) 2.66(0.72) 0.19(0.05) 2.74(0.68) 1.86(0.48) 2.74(0.65)
PVAE-Vote NIPS’24 1.8(0.41) 3.07(0.71) 5.42(1.45) 3.43(0.85) 2.67(0.76) 0.1(0.03) 3.14(0.83) 1.97(0.54) 2.7(0.7)
CP-Vote PRCV’24 3.83(1.06) 3.02(0.66) 6.01(1.05) 4.28(0.92) 2.84(1.25) 0.38(0.15) 4.94(1.24) 2.72(0.88) 3.5(0.9)
GFS arXiv’23 5.13(0.49) 6.13(0.9) 1.91(0.32) 4.39(0.57) 0.15(0.05) 0.02(0.01) 0.48(0.1) 0.22(0.05) 2.3(0.31)
Ours - 3.08(0.49) 3.89(0.49) 16.38(1.06) 7.78(0.68) 1.76(0.22) 2.03(0.64) 11.89(0.96) 5.22(0.61) 6.5(0.64)
Table 13: Performance comparison in mAP(%) on PVRCNN++ detector in the KITTI $\rightarrow 5$shot-A2D2 GCFS task. The bold values represent the best performance. Subscript values in parentheses are standard deviations. The trk for Truck, bcy is for Bicycle, and uvc for Utility_vehicle.
Methods Venues car ped trk common bcy uvc bus novel overall
Target-FT - 3.53(0.91) 2.49(0.58) 4.9(1.52) 3.64(1) 0.31(0.07) 0.07(0.04) 0.41(0.18) 0.26(0.09) 1.95(0.55)
Proto-Vote NIPS’22 1.93(0.43) 1.93(0.39) 7.9(1.6) 3.92(0.81) 1.78(0.29) 0.08(0.04) 2.57(1.16) 1.48(0.5) 2.7(0.65)
PVAE-Vote NIPS’24 1.71(0.83) 2.05(0.54) 8.15(1.5) 3.97(0.96) 1.88(0.51) 0.11(0.08) 2.04(1.02) 1.34(0.54) 2.66(0.75)
CP-Vote PRCV’24 2.69(0.77) 3.14(1.32) 6.77(1.21) 4.2(1.1) 2.49(1.08) 0.06(0.04) 3.87(0.78) 2.14(0.63) 3.17(0.87)
GFS arXiv’23 3.74(0.7) 4.2(0.59) 2.42(0.39) 3.46(0.56) 0.24(0.04) 0.06(0.03) 0.13(0.05) 0.14(0.04) 1.8(0.3)
Ours - 3.8(1.02) 3.52(0.46) 21.94(1.78) 9.75(1.09) 2.88(0.35) 3.14(0.85) 11.28(1.19) 5.76(0.8) 7.76(0.94)
8 Details on Experiments
8.1 Experiment Settings
[…]
Implementation details. For the consistency of input point clouds across datasets, we unify the LiDAR coordinate system of all datasets by setting the origin on the ground. We adopt the point cloud range of [$-75.2m,-75.2m,-2m,75.2m,75.2m,4m$] and the voxel size of [$0.1m,0.1m,0.15m$]. For ground-truth sampling augmentation, we utilize its image-involved version implemented for KITTI in [35] and extend it to A2D2. Regarding class-specific attention module, the head number of is 4 and dropout rate is 0.1. During meta-training, we apply point density-resampling [17] on support data to enlarge domain shifts between query and support data. All experiments are conducted on 2× GeForce RTX-3090. Epoch numbers of the meta-training are limited 5 for NuScenes and Waymo, and 15 for KITTI. Batch sizes are 2 during pre-training and meta-training and 1 in few-shot fine-tuning and testing. $\lambda$, and $\lambda_{1}$, $\lambda_{2}$ are set to 1.0, 0.2, and 0.2.
Table 14: Component ablations in mAP(%) for all classes. (Image-Fusion is our proposed image-guided multi-modal fusion and CL-Proto is our proposed contrastive-learning-enhanced prototype learning.)
Target-
FT
Image-
Fusion
CL-
Proto
car ped trk common van ps cyc trm novel overall
(a) 30.08 7.08 1.17 12.77 14.24 0.87 4.59 2.24 5.48 8.61
(b) 37.07 6.04 1.29 14.80 14.52 0.95 14.40 2.52 8.10 10.97
(c) 35.40 7.21 1.47 14.69 21.16 1.45 19.00 3.08 11.17 12.68
(d) 37.71 7.16 3.11 15.99 22.29 1.54 18.26 4.79 11.72 13.55
Table 15: Performances in mAP(%) with different $K$ for all classes. (Full-shot denotes the detection model trained on all KITTI train data following pre-training settings.)
K car ped trk common van ps cyc trm novel overall
1 19.14 1.62 1.04 7.27 1.88 0.00 0.40 0.00 0.57 3.44
3 27.91 6.28 2.62 12.27 15.62 0.06 9.96 5.42 7.76 9.70
5 37.71 7.16 3.11 15.99 22.29 1.54 18.26 4.79 11.72 13.55
10 55.64 9.38 5.67 23.56 19.04 0.85 26.34 2.60 12.21 17.08
20 56.20 19.38 7.19 27.59 27.92 2.23 33.82 6.01 17.49 21.82
40 62.68 20.40 13.08 32.05 37.89 2.55 39.45 6.30 21.55 26.05
Full-shot 80.64 40.53 2.85 41.34 41.79 0.22 30.68 0.72 18.35 28.21
Compared methods. Indoor FSL methods (i.e., Proto-Vote [53], PVAE-Vote [37], and CP-Vote [18]), are mainly designed for the detection on novel classes, we extend their processing to common classes, to fit the GCFS tasks. Since Proto-Vote [53] is implemented for the indoor RGBD-based data with VoteNet [28] as the base detection model, we extend it to the outdoor LiDAR-abased data with the VoxelRCNN [5] as the base detection model, following our experiment setting. Considering no public codebase for PVAE-Vote and CP-Vote, we follow the paper methodologies and implementation details in papers and extend them to the GCFS tasks. For PVAE-Vote, given that the instability of VAE training is particularly pronounced in outdoor sparse and various point clouds, we incorporate the skip-connection architecture similar to ResNet, which enables VAE branches to learn residuals, thereby enhancing the stability of few-shot training. Regarding outdoor GFSL method GFS-Det with no public codebase, we follow the paper methodology and implementation details in [19] to extend it to the GCFS tasks.
[…]
Table 16: Comparison in mAP(%) with OVD and DA methods under the unsupervised few-shot setting for all classes.
Method Venus car ped trk common van ps cyc trm novel overall
DA SN CVPR’20 19.96 15.11 1.20 12.09 - - - - - -
ST3D++ PAMI’22 56.68 4.66 1.65 21.00 - - - - - -
OVD FnP ECCV’24 20.25 11.11 0.40 10.59 9.19 0.11 0.71 0.62 2.66 6.06
Ours-OVD - 42.67 22.39 1.69 22.25 22.69 1.45 7.77 1.15 8.26 14.26
8.2 Experimental Results
Tables 8, 9, 10, 11, 12, and 13, show more detailed performance among all classes w.r.t. Table 2. As shown in them, across all GCFS tasks, compared to existing methods, our method achieves more accurate object detection performance for overall common and novel classes, especially for “Car”, “Truck”, “Van”, “Cyclist”, “Tram”, “Utility_Vehicle”, and “Bus”. It demonstrates our method’s strong knowledge transferability from the common object in the source domain while effectively generalizing to novel classes with limited few-shot samples. Regarding the evaluation on high-density (64-beam) KITTI or low-density (16-beam) A2D2, the superior performance of our method underscores its strong adaptability to both moderate and extreme domain shifts. Especially in KITTI $\rightarrow 5$shot-A2D2 GCFS task, the performance of our method overpass the second-best performance significantly (i.e., overall mAP: VoxelRCNN $2.74\%\rightarrow 6.5\%$ and PVRCNN++ $3.17\%\rightarrow 7.76\%$), ensuring robust few-shot detection even in challenging low-density scenarios. The results of indoor 3D FSL methods (i.e., Proto-Vote, PVAE-Vote, CP-Vote) reflect the challenges in extending them to outdoor scenarios that is characterized by sparse point clouds at greater distances, dynamic objects, and varying lighting and weather conditions. Especially for common classes shared between source data and target data, those methods struggle with the demands of outdoor environments, resulting in reduced accuracy and robustness in the outdoor detection contexts. GFS-Det performs well mostly in common classes, especially on “Pedestrian” objects, indicating that its dedicated category-specific branches reduce interference of novel objects to common objects well learned in source pre-training. Yet, this separate-branch learning strategy forces the novel-object branch to learn geometric features from scratch, preventing it from leveraging geometric priors from common classes like cars or pedestrians. As a consequence, GFS-Det struggles with novel classes, hindering its ability to generalize effectively to newly learned objects and limiting its adaptability in GCFS tasks.
Table 14 shows the performance of our proposed image-guided multi-modal fusion (denoted as Image-Fusion) method and our proposed contrastive-learning-enhanced prototype learning (denoted as CL-Proto) among all object classes. The experimental results show that our method demonstrates significant advantages in both novel and common classes, especially when combining Image-Fusion and CL-Proto, as in row (d), where it achieves the best performance. Specifically, the introduction of Image-Fusion significantly improves the performance on novel classes, raising the mAP from 5.48% to 11.17%. This improvement is particularly evident in classes like “Van”, “Cyclist”, and “Tram”, where data scarcity makes single-modal features insufficient. Leveraging image-guided multi-modal fusion enables the model to better capture features in novel classes, enhancing adaptability in few-shot scenarios. On the other hand, our proposed contrastive-learning-enhanced prototype learning mainly enhances the performance on common classes. When CL-Proto is added alone, the mAP for common classes increases from 12.77% to 14.80%, with a particularly notable improvement in the “Car” class, where mAP rises from 30.08% to 37.07%. Our proposed contrastive-learning-enhanced prototype learning improves the detection model with intra-class and inter-class differentiation, allowing the model to more accurately identify various features against source and target domain gaps. When Image-Fusion and CL-Proto are combined, as in row (d), the model achieves optimal performance in both novel and common classes, with an overall mAP reaching 13.55%. For novel classes, the mAP increases to 11.72%, and for common classes, it rises to 15.99%. This combination fully leverages the multi-modal feature representation strengths of our proposed image-guided multi-modal fusion method and our proposed contrastive-learning-enhanced prototype learning, enabling the model to perform better in the GCFS task.
Table 15 shows the performance of our method under different $K$-shot target data. The results show that as the number of few-shot samples K increases, the model’s overall performance improves steadily. For instance, when K increases from 1 to 40, the overall mAP rises from 3.44% to 28.21%, indicating that a higher sample count helps the model better learn target features and improve detection accuracy. This trend suggests that with more samples, the model can effectively learn features for categories with abundant data. Regarding the “Van”, “Person_sitting”, and “Tram” categories, performance exhibits irregular fluctuations as $K$ increases (i.e. $5\rightarrow 10$). This variation may stem from the randomness in frame sampling for few-shot conditions. Given the limited frames, the quality of each object can vary, affecting the model’s stability and consistency. Additionally, Full-shot training results indicate that even with training on the entire dataset, certain categories such as “Truck”, “Person_sitting”, “Cyclist”, and “Tram” show relatively low detection accuracy. On one hand, the limited quantity of some categories (488 trucks, 224 trams, and 56 sitting persons w.r.t. 3769 training frames) in the training data restricts the model’s ability to fully learn their features, resulting in lower accuracy. On the other hand, image-guided approach enhances the discovery of novel semantics , boosting recall on novel objects (e.g., “Person_sitting”, “Cyclist”, and “Tram”).
The results in Table 16 show that under the unsupervised few-shot setting, our extended OVD method demonstrates significant advantages across both common and novel classes. Although ST3D++ performs well in the car class, its performance is limited for other common classes, highlighting its lack of generalization in few-shot scenarios. Meanwhile, FnP’s initial advantage is largely due to its cautious box candidate search strategy, which works effectively in traditional OVD settings by leveraging a large amount of target data to accumulate good object samples. However, this approach is evidently inadequate for dealing with few-shot data, due to even fewer object samples for object feature learning. In contrast, our method achieves overall mAPs of 22.25% on common classes and 8.26% on novel classes, with a notable mAP boost on “Pedestrian”, “Van”, and “Cyclist”. This shows that our method, under the unsupervised few-shot setting, can effectively handle feature distribution differences in the target domain, achieving more accurate and balanced detection for both common and novel classes.